OpenStudy (anonymous):

which rule allows you to change the order of the terms when dealing with addition and multiplication which rule allows you to change the order of the terms when dealing with addition and multiplication @Mathematics

OpenStudy (anonymous):

conmutativa

OpenStudy (anonymous):

How would you graph this problem 4x+2z+3x=24

hero (hero):

a + b = b + a ab = ba Is that what you're talking about?